A fuel tank used in a barbeque contains 1.09 x 10-1 moles of propane, C3H8(g). How many molecules of propane are in the tank?
1
Expert's answer
2020-12-10T07:51:56-0500

Solution:

Moles of propane = 1.09×10-1 moles = 0.109 moles

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of propane molecules = 0.109 mol × (6.022×1023 molecules / 1 mol) = 6.56×1022 molecules


Answer: 6.56×1022 molecules of propane are in the tank.